Jack, a student at School of Business is planning to promote a bike rental company in Ohio. The following are the available estimates:

Cost of purchasing bikes = $20,000

Advertising costs = $1000

Permit cost = 500.

Average price per rental = $12.00

Variable cost per rental = $2.50

Calculate the fixed cost, contribution and break even
Answer #1

The fixed cost is the cost that is incurred irrespective of the production level. That is

20000 + 1000 + 500 = 21500

The contribution is the profit per rental. That is

12.00 – 2.50 = 9.50

The breakeven is the fixed cost divided by the contribution. That is

21500/9.50 = 2263.15 or 2264 units rented out